Find the equation of the line that passes through the pair of points. Write your answer in standard form.
(1/4, -1/3) (-1/4, -3)

Answer by cdmcg(2)

use the formula y2-y1=m(x2-x1)
-3+1/3=m(-1/4-1/4)
-8/3=m(-1/2)
m=16/3
standard formula ax+by=c
16/3=(y+3):(x+1/4) => 3(y+3)=16(x+1/4) 3y+9=16x+4 =>-16x+3y=-5 =>16x-3y=5